OpenStudy (anonymous):

Ken is financing a $355,000 mortgage for 30 years at a fixed rate of 7.55%. What is the total cost of the principal and interest after 30 years? $594,178.20 $897,976.80 $763,605.00 $390,668.70

OpenStudy (anonymous):

I=Prt

OpenStudy (anonymous):

it more complex than that

OpenStudy (anonymous):

the rate is fixed therefore it isn't exponential

OpenStudy (anonymous):

ok, can u break it down a little more

jimthompson5910 (jim_thompson5910):

First you need to find the monthly payment P = L[c(1 + c)^n]/[(1 + c)^n - 1] P = 355000[0.00629166666666667(1 + 0.00629166666666667)^360]/[(1 + 0.00629166666666667)^360 - 1] P = 355000[0.00629166666666667(1.00629166666667)^360]/[(1.00629166666667)^360 - 1] P = 355000[0.00629166666666667(9.5630281969118)]/[9.5630281969118 - 1] P = 355000(0.0601673857389034)/(9.5630281969118 - 1) P = 355000(0.0601673857389034)/(8.5630281969118) P = 355000(0.00702641452945377) P = 2494.37715795609 P = 2494.38 So the monthly payment is $2494.38 So over 30 years (or 30*12 = 360 months) you will have paid 2494.38*360 = $897,976.80